OpenStudy (anonymous):

Find the dy/dx for the following functions: i) y = x^2Secx ii) y = 2xe^2x / sqrt x^2 + 3x

OpenStudy (dumbcow):

i) use product rule let f = x^2 .............f' = 2x g = sec(x) ............g' = sec(x)tan(x) --> f'g + fg' = 2xsec(x) + x^2sec(x)tan(x) ii) use quotient rule, then product and chain rules Let f = 2xe^(2x) Find f' : a = 2x .............. a' = 2 b = e^(2x) ..........b' = 2e^(2x) f' = 2e^(2x) + 4xe^(2x) Let g = sqrt(x^2 +3x) Find g': using chain rule, let u = x^2 +3x, du = 2x +3 d/du sqrt(u) = 1/2sqrt(u) g' = (2x+3)/2sqrt(x^2+3x) quotient rule --> (f'g -fg')/g^2 = [ (2e^(2x) + 4xe^(2x))sqrt(x^2 +3x) - 2xe^(2x)(2x+3)/sqrt(x^2 +3x) ] / (x^2+3x) =[(2e^(2x) + 4xe^(2x))(x^2+3x) - 2xe^(2x)(2x+3)] / (x^2+3x)sqrt(x^2 +3x) = [2x^2e^(2x)(2x+5)] / (x^2+3x)sqrt(x^2 +3x)